# Monadic reconstruction of unitary Drinfeld centers and Factorization Homology

### v1

@@ 7. Factorization Homology
We consider now running factorization homology with categories of unitary representations of locally compact quantum groups, and more generally C-tensor categories, as coefficients. For this, we take $\mathcal{A}$ to be the $(2,1)$-category $\operatorname{C*Cat}$: objects are small C-categories, 1-morphisms are *-functors, and 2-morphisms are unitary natural isomorphisms. This category was studied in detail in [MR3007088], where the following is shown.
Theorem 7.4 ([MR3007088], Section 4).
Remark 7.5.
It is then a simple exercise to show that $\mathbb{E}_{2}$-algebras in $\operatorname{C*Cat}$ are small unitarily braided C-tensor categories. With the above discussion, we conclude the following fact.
@@
factorization homology with coefficient in $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$, such that
@@
This gives to $\int_{\Sigma}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$ the structure of a $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$-module C-category.
@@
Suppose $\partial\Sigma\neq\emptyset$. Given a boundary component $S$ of $\Sigma$, consider the induced $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$-module C category structure on $\int_{\Sigma}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$. Then
$\left(\int_{\Sigma}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})\underset{\mathcal{Z}\mathrm{Hilb}(\mathcal{C})}{\boxtimes}\mathrm{Hilb}(\mathcal{C}),\mathcal{O}_{\Sigma}\boxtimes{\mathds{1}}\right)$
has a canonical structure of a centrally pointed $\mathcal{C}$-bimodule C-category. Define
$\mathcal{Y}_{\Sigma}:=\underline{\operatorname{End}}_{\mathcal{C}^{mp}\boxtimes\mathcal{C}}(\mathcal{O}_{\Sigma}\boxtimes{\mathds{1}})\ .$
@@
If $\Sigma$ is a compact framed surface and $S$ is a non-empty boundary component of $\Sigma$, there is an associated C-algebraic extension
$B_{F}\subset(A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{\Sigma}\ .$
If $\Gamma_{S}(\Sigma)$ is the modular group of $\Sigma$ relative to the boundary component $S$, there is a canonical action of $\Gamma_{S}(\Sigma)$ by *-automorphisms on $(A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{\Sigma}$ such that
$B_{F}\subset\left((A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{\Sigma}\right)^{\Gamma_{S}(\Sigma)}\ .$
Moreover, there is a fully-faithful unitary $\mathcal{C}^{mp}\boxtimes\mathcal{C}$-modue functor
$\left(\int_{\Sigma}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})\underset{\mathcal{Z}\mathrm{Hilb}(\mathcal{C})}{\boxtimes}\mathrm{Hilb}(\mathcal{C}),\mathcal{O}_{\Sigma}\boxtimes{\mathds{1}}\right)\to\left(\operatorname{Corr}^{l}((A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{\Sigma},A^{\text{op}}\otimes A),(A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{\Sigma}\right)\ .$
@@ 7.1. Complex Quantum Groups
Let $\mathbb{K}$ be a compact quantum group. Given a compact framed surface $\Sigma$ with a chosen non-empty boundary component $S$, there exists a Yetter-Drinfeld $\mathbb{K}$-C-algebra $A_{\Sigma}$ such that
$\int_{\Sigma}\operatorname{Rep}(D\mathbb{K})\underset{\operatorname{Rep}(D\mathbb{K})}{\boxtimes}\operatorname{Rep}(\mathbb{K})\simeq\mathrm{Hilb}_{A_{\Sigma}}^{\mathbb{K}}\ ,$
as $\operatorname{Rep}(\mathbb{K})$-bimodule C-categories. There is a canonical action of $\Gamma_{S}(\Sigma)$ on $A_{\Sigma}$ by $D\mathbb{K}$-equivariant *-automorphisms.

### v2

@@ 7. Factorization Homology
We consider now running factorization homology with categories of unitary representations of locally compact quantum groups, and more generally C-tensor categories, as coefficients. For this, we take $\mathcal{A}$ to be the $(2,1)$-category $\operatorname{C*Cat}$. It is then a simple exercise to show that $\mathbb{E}_{2}$-algebras in $\operatorname{C*Cat}$ are small unitarily braided C-tensor categories. With the above discussion, we conclude the following fact.
@@
factorization homology with coefficients in $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$, such that
@@
This gives to $\int_{\Sigma}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$ the structure of a $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$-module C-category. More generally, if $\Sigma$ has $n$ boundary components then factorization homology gives a $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})^{\boxtimes n}$-module C-category.
@@
Suppose $\partial\Sigma\neq\emptyset$, with $n$-boundary components. Consider the induced $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})^{\boxtimes n}$-module C-category structure on $\int_{\Sigma}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$. Then
$\left(\int_{\Sigma}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})\underset{\mathcal{Z}\mathrm{Hilb}(\mathcal{C})^{\boxtimes n}}{\boxtimes}\mathrm{Hilb}(\mathcal{C})^{\boxtimes n},\mathcal{O}_{\Sigma}\boxtimes{\mathds{1}}\right)$
has a canonical structure of a centrally pointed $\mathcal{C}^{\boxtimes n}$-bimodule C-category. Define
$\mathcal{Y}_{\Sigma}:=\underline{\operatorname{End}}_{(\mathcal{C}^{mp}\boxtimes\mathcal{C})^{\boxtimes n}}(\mathcal{O}_{\Sigma}\boxtimes{\mathds{1}})\ .$
@@
If $\Sigma$ is a compact framed surface with $n$-boundary components, there is an associated C-algebraic extension
$B_{F}^{\otimes n}\subset(A^{\text{op}}\otimes A)^{\otimes n}\underset{F}{\rtimes}\mathcal{Y}_{\Sigma}\ .$
If $\Gamma(\Sigma)$ is the modular group of $\Sigma$ relative to its boundary, there is a canonical action of $\Gamma(\Sigma)$ by *-automorphisms on $(A^{\text{op}}\otimes A)^{\otimes n}\underset{F}{\rtimes}\mathcal{Y}_{\Sigma}$ such that
$B_{F}^{\otimes n}\subset\left((A^{\text{op}}\otimes A)^{\otimes n}\underset{F}{\rtimes}\mathcal{Y}_{\Sigma}\right)^{\Gamma(\Sigma)}\ .$
Moreover, there is a fully-faithful unitary $(\mathcal{C}^{mp}\boxtimes\mathcal{C})^{\boxtimes n}$-module functor
$\left(\int_{\Sigma}\mathcal{Z}\mathrm{Hilb}(\mathcal{C})\underset{\mathcal{Z}\mathrm{Hilb}(\mathcal{C})^{\boxtimes n}}{\boxtimes}\mathrm{Hilb}(\mathcal{C})^{\boxtimes n},\mathcal{O}_{\Sigma}\boxtimes{\mathds{1}}\right)\to\left(\operatorname{Corr}^{l}((A^{\text{op}}\otimes A)^{\otimes n}\underset{F}{\rtimes}\mathcal{Y}_{\Sigma},(A^{\text{op}}\otimes A)^{\otimes n}),(A^{\text{op}}\otimes A)^{\otimes n}\underset{F}{\rtimes}\mathcal{Y}_{\Sigma}\right)\ .$
@@ 7.1. Complex Quantum Groups
Let $\mathbb{K}$ be a compact quantum group. Given a compact framed surface $\Sigma$ with $n>0$ boundary components, there exists a Yetter-Drinfeld $\mathbb{K}^{\times n}$-C-algebra $A_{\Sigma}$ such that
$\int_{\Sigma}\operatorname{Rep}(D\mathbb{K})\underset{\operatorname{Rep}(D\mathbb{K})^{\boxtimes n}}{\boxtimes}\operatorname{Rep}(\mathbb{K})^{\boxtimes n}\simeq\mathrm{Hilb}_{A_{\Sigma}}^{\mathbb{K}}\ ,$
as $\operatorname{Rep}(\mathbb{K}^{\times n})$-bimodule C-categories. There is a canonical action of $\Gamma(\Sigma)$ on $A_{\Sigma}$ by $D\mathbb{K}^{\times n}$-equivariant *-automorphisms.